# On the Weierstraß form of infinite dimensional differential algebraic equations

### v1

@@ 3 Sufficient condition for a Weierstraß form
We want to show that the radiality index exists by using the Weierstraß form. Define $Q_{b}z=z-\frac{\langle c,z\rangle}{\langle c,b\rangle}b$. Then, $Q_{b}b=0$, ${\rm ran}\,Q_{b}={\rm ker}\,C$ and $\mathcal{W}=\mathcal{W}^{1}\oplus\mathcal{W}^{2}\coloneqq{\rm ker}\,C\oplus\mathrm{span}(b)$. For more concise notation, we additionally define $\tilde{C}=\frac{1}{\langle c,b\rangle}$, $\tilde{B}=\frac{1}{\langle c,b\rangle}B$ and $K\coloneqq\tilde{C}A_{0}$. Then we can define the isomorphisms $U\colon\mathcal{W}\times\mathbb{C}\to\mathcal{W}^{1}\times\mathcal{W}^{2}\times\mathbb{C}$ and $V\colon\mathcal{W}^{1}\times\mathcal{W}^{2}\times\mathbb{C}\to\mathcal{W}\times\mathbb{C}$ via
@@
Next we investigate the radiality index for DAEs in Weierstraß form.
@@
Let $N$ has nilpotency degree $k+1\in\mathbb{N}$ and $\lambda\in\rho(E,A)=\rho(\tilde{E},\tilde{A})$. Then, $\lambda\in\rho(I_{\mathcal{Y}^{1}},A_{1})$. Conversely, if $\lambda\in\rho(I_{\mathcal{Y}^{1}},A_{1})$, then $(\lambda N-I_{\mathcal{Y}^{2}})^{-1}=-\sum_{l=0}^{k}(\lambda N)^{l}$. Thus,
@@
Especially, for $\lambda\in\rho(E,A)$ we obtain
$(\lambda E-A)^{-1}E=\Big{[}\begin{smallmatrix}(\lambda I_{\mathcal{Y}^{1}}-A_{1})^{-1}&0\\ 0&(\lambda N-I_{\mathcal{Y}^{2}})^{-1}N\end{smallmatrix}\Big{]}=\tilde{E}(\lambda\tilde{E}-\tilde{A})^{-1}$ (22)
@@
Thus, Part a) follows directly from from (21) and (23). To show Part b) assume, that $A_{1}$ generates a $C_{0}$-semigroup and $N$ has a nilpotency degree smaller or equal to $p+1$. Using the Theorem of Hille-Yosida there exists $\omega,C>0$ such that $(\omega,\infty)\subseteq\rho(I_{\mathcal{Y}^{1}},A_{1})=\rho(A_{1})$ and $\|((\lambda I_{\mathcal{Y}^{1}}-A_{1})^{-1})^{n}\|\leq\frac{C}{(\lambda-\omega)^{n}}$ for all $\lambda\geq\omega$, $n\in\mathbb{N}$. Together with (22) and (23) we have
@@
[14, eq. (4.1)]. Furthermore, for all $x_{0}\in\mathcal{X}$ the function $t\mapsto S(t)x$ is a solution of the $n$-times integrated Cauchy problem and, thus, a mild solution of (24).
@@
for given $C,\tilde{C}>0$ and $\omega>0$. Thus, $(I_{\mathcal{Y}^{1}},A_{1})$ has a complex resolvent index and by [14, Cor. 4.9] $A_{1}$ generates an (at least) $p+2$-times integrated semigroup. ∎
@@
Assume, that $(E,A)$ has a complex radiality index. Then $A_{1}E_{1}^{-1}$ and $E_{1}^{-1}A_{1}$ from (18w) and (19w) generate integrated semigroups.
@@ 4 Infinite dimensional port-Hamiltonian DAEs
    By b) (26) is equivalent to $\frac{d}{dt}EQ^{-1}z(t)=Az(t)$. Since ${\rm ran}\,E$ is closed, the same holds for ${\rm ran}\,EQ^{-1}$ and one can split up the space into $\mathcal{Z}={\rm ran}\,EQ^{-1}\oplus{\rm ker}\,EQ^{-1}\eqqcolon\mathcal{Z}^{1}\oplus\mathcal{Z}^{2}$. Thus, $\frac{d}{dt}EQ^{-1}z(t)=Az(t)$ is equivalent to
@@
    whereby $A_{i}\colon\mathcal{Z}\to\mathcal{Z}^{1}$ and $E_{1}\coloneqq EQ^{-1}_{{\rm ran}\,EQ^{-1}}$. Thus, the representation for port-Hamiltonian differential algebraic equations chosen here corresponds with the notion of an abstract Hamiltonian differential-algebraic equation defined in [12, Assumption 7]. Furthermore, it should be mentioned that there are more ways to define and approach port-Hamiltonian systems, such as using relations [3] or Dirac-structures [23].
@@
Since ${\rm ran}\,B$ is closed and $B$ is self-adjoint, one can decompose $\mathcal{Z}$ into $\mathcal{Z}={\rm ran}\,B\oplus{\rm ker}\,B$. Let $P_{{\rm ran}\,B}$ and $P_{{\rm ker}\,B}$ be the projections of $\mathcal{Z}$ onto ${\rm ran}\,B$ and ${\rm ker}\,B$. Let $z=P_{{\rm ran}\,B}z+P_{{\rm ker}\,B}z\in\mathcal{Z}$ be arbitrary. Then there exists a $y\in\mathcal{Z}$ with $P_{{\rm ran}\,B}z=By$. Define
$Tz\coloneqq P_{{\rm ran}\,B}y+P_{{\rm ker}\,B}z.$
Then $T$ is indeed well defined. To be more precise, choose $y_{1},y_{2}\in\mathcal{Z}$ and assume $By_{1}=By_{2}=P_{{\rm ran}\,B}z$. Thus $y_{1}-y_{2}\in{\rm ker}\,B$ and, therefore, $P_{{\rm ran}\,B}(y_{1}-y_{2})=0$. By simple calculations one has
$BTBz=B(P_{{\rm ran}\,B}z)=B(P_{{\rm ran}\,B}z)+B(P_{{\rm ker}\,B}z)=Bz.$
What remains to be shown is that there exists a $c>0$, such that $T\geq cI_{\mathcal{Z}}$. Since $B$ is non-negative and self-adjoint there exists a non-negative and self-adjoint operator $V\in L(\mathcal{Z})$ with $B=V^{2}$. Thus, ${\rm ran}\,B\subseteq{\rm ran}\,V$ and by [24, Thm. 2.2 (p.208)] there exists a $\varepsilon>0$ with
$\left\|Bz\right\|_{\mathcal{Z}}\leq\varepsilon\left\|Vz\right\|_{\mathcal{Z}}$
holds for all $z\in\mathcal{Z}$. Let $x,y\in\mathcal{Z}$ with $By=P_{{\rm ran}\,B}z$. Hence,
$\langle z,Tz\rangle_{\mathcal{Z}}$ $=\langle P_{{\rm ran}\,B}z+P_{{\rm ker}\,B}z,P_{{\rm ran}\,B}y+P_{{\rm ker}\,B}z\rangle_{\mathcal{Z}}$
$=\langle P_{{\rm ran}\,B}z,P_{{\rm ran}\,B}y\rangle_{\mathcal{Z}}+\langle P_{{\rm ker}\,B}z,P_{{\rm ker}\,B}z\rangle_{\mathcal{Z}}$
$=\langle By,P_{{\rm ran}\,B}y\rangle_{\mathcal{Z}}+\left\|P_{{\rm ker}\,B}\right\|_{\mathcal{Z}}^{2}$
$=\langle y,\underbrace{BP_{{\rm ran}\,B}y}_{=V^{2}y}\rangle_{\mathcal{Z}}+\left\|P_{{\rm ker}\,B}\right\|_{\mathcal{Z}}^{2}$
$=\langle Vy,Vy\rangle_{\mathcal{Z}}+\left\|P_{{\rm ker}\,B}\right\|_{\mathcal{Z}}^{2}$
$\geq\frac{1}{\varepsilon}\langle By,By\rangle_{\mathcal{Z}}+\left\|P_{{\rm ker}\,B}\right\|_{\mathcal{Z}}^{2}=\frac{1}{\varepsilon}\left\|P_{{\rm ran}\,B}\right\|_{\mathcal{Z}}^{2}+\left\|P_{{\rm ker}\,B}\right\|_{\mathcal{Z}}^{2}.$
Thus, for $c\coloneqq\min\{\frac{1}{c},1\}$ one has $\langle z,Tz\rangle_{\mathcal{Z}}\geq c\langle z,z\rangle_{\mathcal{Z}}$. ∎
@@
$\frac{d}{dt}\langle Ex(t),Qx(t)\rangle_{\mathcal{Z}}$ $=\frac{d}{dt}\langle x(t),\underbrace{E^{\ast}Q}_{E^{\ast}TE}x(t)\rangle_{\mathcal{Z}}$
@@
$=2{\rm Re}\langle Qx(t),AQx(t)\rangle_{\mathcal{Z}}\leq 0,\quad t\geq 0.$
Note, that we used the continuity of $E$, $Q$ and $T$ in the second and third equation. ∎
@@
It is possible to show the same for the adjoint system and, if the set containing all trajectories of the solutions of (26) is closed, one can apply [6, Thm. 3.7] to gain a Weierstraß form on a subset. To be more precise, we know that ${\rm ran}\,E^{\ast}$ is closed and $Q^{-1}$ is invertible. Thus, there also exists a $S\in L(\mathcal{X})$ with $S\geq\tilde{c}I_{\mathcal{X}}$, for a $\tilde{c}>0$, with
@@
Thus, if $U\coloneqq\{u\in\mathcal{Z}\,|\,\exists x$ classical solution of (26) $\exists t\geq 0\colon x(t)=u\}\subseteq\mathcal{Z}$ and $V\coloneqq\{v\in\mathcal{Z}\,|\,\exists z$ classical solution of (33) $\exists t\geq 0\colon z(t)=v\}\subseteq\mathcal{Z}$ are both closed sets, one can apply [6, Thm. 3.7] to $(E|_{U},(AQ)|_{U})$ on the spaces $\mathcal{X}_{S^{-1}}$ and $\mathcal{Z}_{T}$.
@@
Consider longitudinal vibrations in a viscoelastic nanorod. Let $l$ be the length of the nanorod, $N(x,t)$ be the resultant force of axial stress, $w(x,t)$ be the displacement of the nanorod in $x$ direction, $C$ be the elastic modulus, $D$ be the cross sectional area, $\mu$ be a nonlocal parameter, $\rho$ the mass density, $\tau_{d}$ the viscous damping and $a^{2}$, $b^{2}$ be the stiffness and damping coefficients of the light viscoelastic layer and consider the system introduced in [8].
@@
In [5] the existence of solution was studied by reducing the system to a homogeneous port-Hamiltonian system (see [7, Ch. 7]). Here, we provide another approach by simply examining the dissipativity of the DAE as seen in Remark 4.5.
@@
Now, as in [5, p. 452] we impose the boundary conditions directly on the space and show, that (36) has radiality index 00. Define
@@
with the inner product $\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}_{t}}\coloneqq\langle Q\cdot,\cdot\rangle_{\mathcal{X}_{t}}$. Since $Q$ is coercive this induces a norm equivalent to $\|\cdot\|_{\mathcal{X}}$. $\mathcal{X}_{t}$ is in fact closed [5, Lem. 4.1]. Define $\tilde{\mathcal{X}}\coloneqq(\mathcal{X},\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}_{t}})$. Then,
@@
holds for all $z=(z_{1},\ldots,z_{5})\in\mathcal{D}(AQ)$. Furthermore, $(AQ)^{\ast}=A^{\ast}Q$ in $\mathcal{X}_{t}$ and, through similar calculations, one derives ${\rm Re}\langle A^{\ast}Qz,Ez\rangle_{\mathcal{X}_{t}}\leq 0$ for all $z\in\mathcal{D}((AQ)^{\ast})$. Then, by [6, Thm. 3.6, 3.7] $(E,AQ)$ is 00-radial and admits a decomposition as in (18w) on $\mathcal{X}_{t}=\mathcal{X}_{t}^{1}\oplus\mathcal{X}_{t}^{2}$ with $E_{2}(AQ)_{2}^{-1}=0$ and $(AQ)_{2}^{-1}E_{2}=0$. In particular, $(AQ)_{1}E_{1}^{-1}$ generates a contraction semigroup on $\mathcal{X}_{t}^{1}$.
@@
By [6, Thm. 3.7] it becomes clear that if $(E,AQ)$ is a pH-DAE with ${\rm Re}\langle AQx,Ex\rangle\leq 0$ and ${\rm Re}\langle(AQ)^{\ast}z,E^{\ast}z\rangle\leq 0$, then $(E,AQ)$ is in fact 00-radial (and therefore has radiality index 00), admits a Weierstraß form as in (18w) or (19w) and $(AQ)_{1}E_{1}^{-1}$ and $E_{1}^{-1}(AQ)_{1}$ generate contraction semigroups on $\tilde{\mathcal{Z}}^{1}$ and $\tilde{\mathcal{X}}^{1}$. In this case, $(AQ)_{1}E_{1}^{-1}$ and $E_{1}^{-1}(AQ)_{1}$ also generate $C_{0}$-semigroups on $\mathcal{Z}^{1}$ and $\mathcal{X}^{1}$.

### v2

@@ 3 Sufficient condition for a Weierstraß form
We want to show that the radiality index exists by using the Weierstraß form. Define $Q_{b}z=z-\frac{\langle c,z\rangle}{\langle c,b\rangle}b$. Then, $Q_{b}b=0$, ${\rm ran}\,Q_{b}={\rm ker}\,C$ and $\mathcal{W}=\mathcal{W}^{1}\oplus\mathcal{W}^{2}\coloneqq{\rm ker}\,C\oplus\mathrm{span}(b)$. For more concise notation, we additionally define $\tilde{C}=\frac{1}{\langle c,b\rangle}C$, $\tilde{B}=\frac{1}{\langle c,b\rangle}B$ and $K\coloneqq\tilde{C}A_{0}$. Note that $Kz\coloneqq\tilde{C}A_{0}z=\frac{1}{\langle c,b\rangle}\langle c,A_{0}z\rangle=\frac{1}{\langle c,b\rangle}\langle A_{0}^{\ast}c,z\rangle$ holds for all $z\in W$ and therefore $K$ is a bounded operator on $\mathcal{W}$. Thus, we can define the isomorphisms $U\colon\mathcal{W}\times\mathbb{C}\to\mathcal{W}^{1}\times\mathcal{W}^{2}\times\mathbb{C}$ and $V\colon\mathcal{W}^{1}\times\mathcal{W}^{2}\times\mathbb{C}\to\mathcal{W}\times\mathbb{C}$ via
@@
Further examples of DAEs with existing radiality index (such as the linearized Navier-stokes equation) can be found in [3]. Next, the radiality index for DAEs in Weierstraß form gets investigated.
@@
Let $N$ has nilpotency degree $k+1\in\mathbb{N}$ and $\lambda\in\rho(E,A)$. Then, $\lambda\in\rho(I_{\mathcal{Y}^{1}},A_{1})$. Conversely, if $\lambda\in\rho(I_{\mathcal{Y}^{1}},A_{1})$, then $(\lambda N-I_{\mathcal{Y}^{2}})^{-1}=-\sum_{l=0}^{k}(\lambda N)^{l}$. Thus,
@@
In particular, for $\lambda\in\rho(E,A)$ we obtain
$(\lambda E-A)^{-1}E=\begin{bmatrix}(\lambda I_{\mathcal{Y}^{1}}-A_{1})^{-1}&0\\ 0&(\lambda N-I_{\mathcal{Y}^{2}})^{-1}N\end{bmatrix}=E(\lambda E-A)^{-1}$ (24)
@@
Thus, Part a) follows directly from (23) and (25). To show Part b) assume, that $A_{1}$ generates a $C_{0}$-semigroup and $N$ has a nilpotency degree smaller or equal to $p+1$. Using the Theorem of Hille-Yosida there exists $\omega,C>0$ such that $(\omega,\infty)\subseteq\rho(I_{\mathcal{Y}^{1}},A_{1})=\rho(A_{1})$ and $\|((\lambda I_{\mathcal{Y}^{1}}-A_{1})^{-1})^{n}\|\leq\frac{C}{(\lambda-\omega)^{n}}$ for all $\lambda\geq\omega$, $n\in\mathbb{N}$. Together with (24) and (25) we have
@@
[15, eq. (4.1)]. Furthermore, for all $x_{0}\in\mathcal{X}$ the function $t\mapsto S(t)x$ is a solution of the $n$-times integrated Cauchy problem and, thus, a mild solution of (26).
@@
for given $C,\tilde{C}>0$ and $\omega>0$. Thus, $(I_{\mathcal{Y}^{1}},A_{1})$ has a complex resolvent index and by [15, Cor. 4.9] $A_{1}$ generates an (at least) $p+2$-times integrated semigroup. ∎
@@
Assume, that $(E,A)$ has a complex radiality index. Then $A_{1}E_{1}^{-1}$ and $E_{1}^{-1}A_{1}$ from (20) and (21) generate integrated semigroups.
@@ 4 Infinite dimensional port-Hamiltonian DAEs
    By c) (28) is equivalent to $\frac{d}{dt}EQ^{-1}z(t)=Az(t)$. Since ${\rm ran}\,E$ is closed, the same holds for ${\rm ran}\,EQ^{-1}$ and one can split up the space into $\mathcal{Z}={\rm ran}\,EQ^{-1}\oplus{\rm ker}\,EQ^{-1}\eqqcolon\mathcal{Z}^{1}\oplus\mathcal{Z}^{2}$. Thus, $\frac{d}{dt}EQ^{-1}z(t)=Az(t)$ is equivalent to
@@
    whereby $A_{i}\colon\mathcal{Z}\to\mathcal{Z}^{1}$ and $E_{1}\coloneqq EQ^{-1}_{{\rm ran}\,EQ^{-1}}$. Thus, the representation for port-Hamiltonian differential algebraic equations chosen here corresponds with the notion of an abstract Hamiltonian differential-algebraic equation defined in [13, Assumption 7]. Furthermore, it should be mentioned that there are more ways to define and approach port-Hamiltonian systems, such as using relations [4] or Dirac-structures [24].
@@
Since ${\rm ran}\,B$ is closed and $B$ is self-adjoint, $\mathcal{Z}$ has an orthogonal decomposition into $\mathcal{Z}={\rm ran}\,B\,\oplus\,{\rm ker}\,B$. Let $P_{{\rm ran}\,B}$ and $P_{{\rm ker}\,B}$ be the projections of $\mathcal{Z}$ onto ${\rm ran}\,B$ and ${\rm ker}\,B$. Define $\tilde{B}\colon{\rm ran}\,B\to{\rm ran}\,B$, $\tilde{B}(z)=Bz$. This is a bounded, positive selfadjoint operator on the Hilbert space ${\rm ran}\,B$, which is bijective (note that ${\rm ker}\,\tilde{B}={\rm ker}\,B\cap{\rm ran}\,B=\{0\}$). Hence, since $\tilde{B}^{-1}$ is bounded, $\tilde{B}$ is strictly positive with
$\langle\tilde{B}z_{1},z_{1}\rangle\geq c\|z_{1}\|^{2},\quad z_{1}\in{\rm ran}\,B$
for all $z_{1}\in{\rm ran}\,B$, for some $c>0$.
Let $\iota_{{\rm ran}\,B}\colon{\rm ran}\,B\to\mathcal{Z}$ be the embedding and define
$T\colon\iota_{{\rm ran}\,B}\tilde{B}^{-1}\iota_{{\rm ran}\,B}^{\ast}+P_{{\rm ker}\,B}.$
This operator satisfies $BTB=B$ and is bounded. Let $z=z_{1}+z_{2}\in\mathcal{Z}={\rm ran}\,B\,\oplus{\rm ker}\,B$. Then
$\langle Tz,z\rangle$ $=\langle\tilde{B}z_{1}+z_{2},z_{1}+z_{2}\rangle$
$=\langle\tilde{B}z_{1},z_{1}\rangle+\langle z_{2},z_{2}\rangle$
$\geq c\|z_{1}\|^{2}+\|z_{2}\|^{2}.$
Thus, $T$ is strictly positive. The self-adjointness of $T$ follows from that of $B$, completing the proof. ∎
@@
$\frac{d}{dt}\langle Ex(t),Qx(t)\rangle_{\mathcal{Z}}$ $=\frac{d}{dt}\langle x(t),\underbrace{E^{\ast}Q}_{=E^{\ast}TE}x(t)\rangle_{\mathcal{Z}}$
@@
$=2{\rm Re}\langle Qx(t),AQx(t)\rangle_{\mathcal{Z}}\leq 0,\quad t\geq 0,$
where we used the continuity of $E$, $Q$ and $T$ in the second and third equation. Note, that we used $\langle Ex(t)=TEx(t)\rangle=\langle x(t),E^{\ast}TEx(t)\rangle=\langle x(t),Q^{\ast}Ex(t)\rangle=\langle Qx(t),Ex(t)\rangle$ in the third equality. ∎
@@
It is possible to show the same for the adjoint system and, if the set containing all trajectories of the solutions of (28) is closed, one can apply [7, Thm. 3.7] to gain a Weierstraß form on a subset. To be more precise, we know that ${\rm ran}\,E^{\ast}$ is closed and $Q^{-1}$ is invertible. Thus, there also exists a $S\in L(\mathcal{X})$ with $S\geq\tilde{c}I_{\mathcal{X}}$, for a $\tilde{c}>0$, with
@@
Thus, if $U\coloneqq\{u\in\mathcal{Z}\,|\,\exists x$ classical solution of (28) $\exists t\geq 0\colon x(t)=u\}\subseteq\mathcal{Z}$ and $V\coloneqq\{v\in\mathcal{Z}\,|\,\exists z$ classical solution of (35) $\exists t\geq 0\colon z(t)=v\}\subseteq\mathcal{Z}$ are both closed sets, one can apply [7, Thm. 3.7] to $(E|_{U},(AQ)|_{U})$ on the spaces $\mathcal{X}_{S^{-1}}$ and $\mathcal{Z}_{T}$.
@@
Consider longitudinal vibrations in a viscoelastic nanorod. Let $l$ be the length of the nanorod, $N(x,t)$ be the resultant force of axial stress, $w(x,t)$ be the displacement of the nanorod in $x$ direction, $C$ be the elastic modulus, $D$ be the cross sectional area, $\mu$ be a nonlocal parameter, $\rho$ the mass density, $\tau_{d}$ the viscous damping and $a^{2}$, $b^{2}$ be the stiffness and damping coefficients of the light viscoelastic layer and consider the system introduced in [9]. Consider
@@
In [6] the existence of solution was studied by reducing the system to a homogeneous port-Hamiltonian system (see [8, Ch. 7]). Here, we provide another approach by simply examining the dissipativity of the DAE as seen in Remark 4.5.
@@
Now, as in [6, p. 452] we impose the boundary conditions directly on the space and show, that (38) has radiality index 00. Define
@@
with the inner product $\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}_{t}}\coloneqq\langle Q\cdot,\cdot\rangle_{\mathcal{X}_{t}}$. Since $Q$ is coercive this induces a norm equivalent to $\|\cdot\|_{\mathcal{X}}$. $\mathcal{X}_{t}$ is in fact closed [6, Lem. 4.1]. Define $\tilde{\mathcal{X}}\coloneqq(\mathcal{X},\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}_{t}})$. Then,
@@
holds for all $z=(z_{1},\ldots,z_{5})\in\mathcal{D}(AQ)$. Furthermore, $(AQ)^{\ast}=A^{\ast}Q$ in $\mathcal{X}_{t}$ and, through similar calculations, one derives ${\rm Re}\langle A^{\ast}Qz,Ez\rangle_{\mathcal{X}_{t}}\leq 0$ for all $z\in\mathcal{D}((AQ)^{\ast})$. Then, by [7, Thm. 3.6, 3.7] $(E,AQ)$ is 00-radial and admits a decomposition as in (20) on $\mathcal{X}_{t}=\mathcal{X}_{t}^{1}\oplus\mathcal{X}_{t}^{2}$ with $E_{2}(AQ)_{2}^{-1}=0$ and $(AQ)_{2}^{-1}E_{2}=0$. In particular, $(AQ)_{1}E_{1}^{-1}$ generates a contraction semigroup on $\mathcal{X}_{t}^{1}$.
@@
By [7, Thm. 3.7] it becomes clear that if $(E,AQ)$ is a pH-DAE with ${\rm Re}\langle AQx,Ex\rangle\leq 0$ and ${\rm Re}\langle(AQ)^{\ast}z,E^{\ast}z\rangle\leq 0$, then $(E,AQ)$ is in fact 00-radial (and therefore has radiality index 00), admits a Weierstraß form as in (20) or (21) and $(AQ)_{1}E_{1}^{-1}$ and $E_{1}^{-1}(AQ)_{1}$ generate contraction semigroups on $\tilde{\mathcal{Z}}^{1}$ and $\tilde{\mathcal{X}}^{1}$. In this case, $(AQ)_{1}E_{1}^{-1}$ and $E_{1}^{-1}(AQ)_{1}$ also generate $C_{0}$-semigroups on $\mathcal{Z}^{1}$ and $\mathcal{X}^{1}$.